OpenStudy (anonymous):

How does the graph of f(x)=3(4)^x-5+2/3 relate to its parent function? I am struggling with this question and honestly don't know how to do it. Any input on the possible answers? *Please select all correct answers* A- The parent function has been translated to the right. B- The parent function has been compressed. C- The parent function has been translated up. D- The parent function has been stretched.

OpenStudy (anonymous):

Well let's try and establish what the parent function is here.

OpenStudy (anonymous):

If I gave you (x-2)^2 - 3, the parent function is the parabolic function, or x^2. That being said, what do you think the parent function of your problem is?

OpenStudy (anonymous):

\[f(x)=3(4)^{x-5}+2/3\]

OpenStudy (anonymous):

I don't know what a parent function is. I honestly have no clue where to start this problem, or how to approach it.

OpenStudy (anonymous):

A parent function, as I will define it, is the most basic function by which another is based upon. A better way to think of this is to see that the fundamental functions in algebra are all parent functions. These are some of them. x x^2 x^3 sqrt(x) abs(x)

OpenStudy (anonymous):

So when we look at any variation of the (let's say the x function), we know that 'x' is the parent function. Ex. 1 f(x) = x-1 relates to it's parent function 'x' by a vertical shift of -1 Ex. 2 g(x) = (x-3)^2 + 1 relates to it's parent function 'x^2' by a vertical shift of +1 and a horizontal shift of +3 (x-shifts are always opposite of the sign).

OpenStudy (anonymous):

f(x) = x + 1 g(x) = 2x + 3 h (x) = x - 5 These are all based on the parent function: f(x) = x

OpenStudy (anonymous):

Am I confusing you?

OpenStudy (mathmath333):

u can simply graph the function here an see the changes

OpenStudy (mathmath333):

 this is graph for \(\large\tt \begin{align} \color{black}{y=4^x}\end{align}\)

OpenStudy (mathmath333):

 and this is graph for \(\large\tt \begin{align} \color{black}{y=3(4)^{x-5}+2/3}\end{align}\)

OpenStudy (mathmath333):

can u see the diffference